40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 13 03:58:27 2015, Jenne, Update, LSC, Transitioned to ASDC MICH (PRMI and PRFPMI) 
    Reply  Tue Jan 13 15:11:30 2015, diego, Update, LSC, Transitioned to ASDC MICH (PRMI and PRFPMI) 
Message ID: 10891     Entry time: Tue Jan 13 03:58:27 2015     Reply to this: 10896
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Transitioned to ASDC MICH (PRMI and PRFPMI) 

[Jenne, Diego, EricQ]

Hopefully there will be more later, but Chiara just went down (network? other?  Q is in there right now looking at it), so this is a so-far-tonight elog.

We have successfully transitioned MICH over from REFL33Q to ASDC in both the PRMI and PRFPMI configurations.  Next up is to start reducing the CARM offset.


Resetting the REFL demod phases

I have been unable to lock the PRMI for more than teeny blips since Thursday.  So, tonight I finally got it to lock with MICH on AS55Q and PRCL on REFL33I, and used that to set the demod phases. 

PRMI, AS55Q and REFL33I MICH PRCL
Input matrix 1*AS55Q 1*R33I
Gain -7 -0.022
DoF trigger POP22I; 80up, 2down POP22I; 80up, 2down
FM trigger FM 2,6,8; 100up, 2down; 0.3sec FM 2,6,9; 100up, 2down; 0.01sec
Normalization n/a n/a
Output matrix 0.5*BS, -0.2625*PRM 1*PRM
UGF servo n/a

n/a

Setting the demod phases, I used an oscillation of 100 cts to PRM, at 400.123 Hz.

REFL 33 demod phase started at 148deg, now 133.2deg.

REFL165 phase started at -105.53, now -172.

No signal in REFL55????  Time series and spectra both look just like noise.  Need to check alignment of beam on PD, or if cables unplugged!!

REFL11 phase started at 16.75, now 18.9deg.

Was then able to lock on REFL 33 I&Q, like normal.

PRMI, REFL33I&Q MICH PRCL
Input matrix 1*R33Q 1*R33I
Gain 2.5 -0.02
DoF trigger POP22I; 80up, 2down POP22I; 80up, 2down
FMs FM 4,5 on FM 4,5 on
FM trigger FM 2,6,8; 100up, 2down; 0.3sec FM 2,6,9; 100up, 2down; 0.01sec
Normalization n/a n/a
Output matrix 0.5*BS, -0.2625*PRM 1*PRM
UGF servo n/a

n/a



Transitioning PRMI from REFL33Q to ASDC

With the PRMI locked on REFL33 I&Q, I found that a MICH offset of -5 counts gives a minimum in ASDC.  From my earlier elog this evening (http://nodus.ligo.caltech.edu:8080/40m/10887), I expect the minimum to be at +1.4nm.  This is only one point though, so I don't know the calibration of the MICH offset yet (we should get this calib during the day by looking at MICH-only).  Anyhow, this informed which side was positive and negative relative to my Optickle plots, so I know that I wanted positive offset in the MICH servo.

I was able to comfortably hold lock at +20 counts.  Looking at a calibration line at 143.125 Hz, I determined that I wanted the matrix element for ASDC to be -0.05.  After I made that transition using ezcastep, I put the POPDC normalization in.  At the time, POPDC was about 151counts, so I put 1/151 in the POPDC->Mich matrix element.

So, here were the final lock parameters.  Note that in PRMI-only, you can acquire lock like this, and with a variety of MICH offsets:

PRMI, ASDC and REFL33I MICH PRCL
Input matrix -0.05 * ASDC 1*R33I
Gain 2.5 -0.02
DoF trigger POP22I; 80up, 2down POP22I; 80up, 2down
FMs FM 4,5 on FM 4,5 on
FM trigger FM 2,6,8; 100up, 2down; 0.3sec FM 2,6,9; 100up, 2down; 0.01sec
Normalization  0.0066 n/a
Output matrix 0.5*BS, -0.2625*PRM 1*PRM
UGF servo n/a

n/a


Locking PRMI part of PRFPMI

Since the PRMI has been fussy, I'm including a brief note on the PRMI settings when the arms are held with ALS off by roughly 3nm.  To get to this point, we just ran the usual carm_cm_up script, and didn't let it run anymore when it asked for confirmation that PRMI was locked.

PRFPMI, Arms=ALS, PRMI=REFL33I&Q MICH PRCL CARM DARM
Input matrix 1*R33Q 1*R33I -1*alsX+1*alsY 1*alsX+1*alsY
Gain 2.5 -0.012 8 8
DoF trigger POP22I; 80up, 2down POP22I; 80up, 2down n/a n/a
FMs FM 4,5 on FM 4,5 on FM 1,2,3,5,6 FM 1,2,3,5,6
FM trigger FM 2,6,8; 50up, 2down; 0.3sec FM 1,2,6,9; 50up, 2down; 0.01sec n/a n/a
Normalization n/a n/a n/a n/a
Output matrix 0.5*BS, -0.2625*PRM 1*PRM 1*etmx+1*etmy -1*etmx+1*etmy
UGF servo n/a set to 150Hz n/a n/a

With MICH offset of -30 counts, AS port is pretty bright.  ASDC dark offset is set to -475.4 by the LSCoffsets script. with MICH offset = 0, ASDC_OUT is around 300counts.  But, with MICH offset = -30, ASDC_OUT is about 525 counts.  So, I put that 525 counts into the ASDC filterbank offset (so it is now the dark offset + this extra offset), so the ASDC offset is currently around -1,000.  This makes the ASDC signal roughly zero when I am ready to transition MICH over to it.  In principle I should probably set it so the average is the same as the MICH offset, but the noise is so high relative to that offset, that it doesn't matter.

After this, we engaged the CARM and DARM UGF servos.  MICH was gain peaking, so I think we might want to turn that one on too, rather than my by-hand turning down the gain.

The transition has been successful 4 or 5 times with the arms held off resonance at 3nm.  Once, we reduced the CARM offset as low as 1.7 (and had to lower the MICH gain to 1.5), but we were still hearing a woomp-woomp sound.  Not sure what that was from.  At this point, Chiara died, so we lost lock.  After that, we re-acquired lock a few more times, but MC keeps losing it.  We are still able to make the MICH to ASDC transition though, which is good.

The transition won't work if the PRCL UGF servo is not on.  The gain multiplier goes from about 1.1 up to 2.4, so the PRCL gain is certainly changing through the transition.

Diego has written up scripts for the individual UGF servos (look for an elog from him separately), so now the carm_cm_up script goes as far as locking the PRMI on REFL33 I&Q, and then it starts to transition.  PRCL UGF is engaged, MICH offset is set to -30 counts, MICH is transitioned to ASDC, POP normalization engaged, CARM UGF servo turned on, and DARM UGF servo turned on.  There are "read"s in the script before each step, so you can stop whenever you like.

Here's the final configuration for the PRFPMI while the arms are held at 3nm, with MICH on ASDC (so after the transition):

PRFPMI, Arms=ALS, PRCL=REFL33I, MICH=ASDC MICH PRCL CARM DARM
Input matrix 0.27*ASDC 1*R33I -1*alsX+1*alsY 1*alsX+1*alsY
Gain 2.5 -0.015 8 8
DoF trigger POP22I; 80up, 2down POP22I; 80up, 2down n/a n/a
FMs FM 4,5 on FM 4,5 on FM 1,2,3,5,6 FM 1,2,3,5,6
FM trigger FM 2,6,8; 50up, 2down; 0.3sec FM 1,2,6,8,9; 50up, 2down; 0.01sec n/a n/a
Normalization 0.0042*POPDC n/a n/a n/a
Output matrix 0.5*BS, -0.2625*PRM 1*PRM 1*etmx+1*etmy -1*etmx+1*etmy
UGF servo n/a set to 150.001Hz set to 115.1Hz set to 110.1

 

The transition for MICH to ASDC has been successful with the arms held off resonance several times tonight. It's all part of the carm_cm_up script now. I think that if we hadn't lost about an hour of time and our momentum, we would have gotten farther.  I have high hopes for tomorrow night!

ELOG V3.1.3-